ghvhjvgyjftyfvgh:

Point M is between points N and O on 2004-01-01-02-00_files/i0210000.jpg. Find the length of 2004-01-01-02-00_files/i0210001.jpg if MO = 12.3 and NO = 26.9.

AZ:

 I can only assume they're asking for the length of NM

AZ:

The length of the whole line NO is 26.9 and you split it into 2 segments NM and MO. And the length of MO is 12.3. This is the relationship we get: NO = NM + MO Plug in the numbers they've already given you 26.9 = NM + 12.3 Can you solve for NM? Imagine that it's x 26.9 = x + 12.3 To solve for x, all you have to do is subtract 12.3 on both sides. What do you get?